# HOD-informed prior for EFT-based full-shape analyses of LSS

### v1

@@ 1 Introduction
In this paper, we use a sample of such scenarios to derive an HOD-informed prior for the EFTofLSS model to address the issue of an overly large nuisance parameter volume. Our method involves generating a diverse set of galaxy power spectrum measurements from HOD mock catalogs, covering a wide range of HOD parameter space and various cosmological models. By fitting the EFTofLSS model to the power spectra derived from these mock catalogs, we identify the best-fit EFTofLSS parameters and use their distribution to form our HOD-informed prior. This approach ensures that the HOD-informed prior encompasses most of the physically plausible power spectra, allowing us to delineate a more constrained and physical parameter space. Regions outside this prior are likely unphysical and can be excluded from the analysis, thereby helping to reduce the prior effect and improve the precision of cosmological parameter estimates. We carefully tested several configurations that could influence the prior, including the impact of different cosmologies, variations in HOD parameter samples, the inclusion or exclusion of the hexadecapole moment, and dependence on covariance matrices. These tests ensure the robustness and reliability of the HOD-informed prior across various scenarios. [59, 60] applied a similar strategy, by incorporating simulation-based priors, in their full-shape analysis, and achieving a reduction of prior volume of bias parameters. However, they were limited in the range of cosmologies they could consider when forming the prior, leading to concerns that the prior might be biased towards a particular set of cosmological parameters. By using a novel emulator for the EFTofLSS we are able to consider a far wider range of cosmologies making this technique feasible for full fits.
@@ 2.2 Galaxy power spectrum in the EFTofLSS
Our notation follows that established in [17], incorporating a combination of linear terms, 1-loop Standard Perturbation Theory (SPT) terms, counterterms, and stochastic terms. The quantity $\mu$ is the cosine of the angle between the LoS direction and wavenumber vector $\mathbf{k}$. $P_{11}(k)$ denotes the linear matter power spectrum, and f is the growth factor. The scale $k_{\textsc{m}}^{-1}$ controls the spatial extension of the collapsed objects, the scale $k_{\textsc{r}}^{-1}$ controls the counterterms needed to define the product of velocities at the same point [77, 78], and $\bar{n}_{g}$ is the mean galaxy number density. $Z_{n}$ represents the $n$th order of the redshift galaxy density kernel,
@@ 3.1 Generating galaxy power spectra from HOD and N-body simulations
We use the AbacusSummit simulation suite [82, 83, 84, 85, 86] for this analysis, which provides a set of large, high-resolution N-body simulations specifically designed to meet the accuracy requirements of the DESI galaxy survey. These simulations cover a wide range of cosmological parameters, making them particularly suitable for our study. For this work, we utilize the AbacusSummit_base_c000_ph000 box, which represents the Planck 2018 cosmology [34], along with the AbacusSummit_base_c130_ph000 to AbacusSummit_base_c160_ph000 boxes at $z=0.8$. The latter form an unstructured emulator grid surrounding the c000 box, providing wider coverage of the 7-dimensional cosmological space $\{\omega_{b}h^{2},\omega_{c}h^{2},h,A_{s},n_{s},w_{0},w_{a}\}$. Each simulation box has a size of $2000\,h^{-1}\rm{Mpc}$ per side and contains $6912^{3}$ dark matter particles. This setup allows us to generate galaxy power spectra with diverse cosmological parameters, ensuring generality in our findings - the prior should not be tied to a single comsological model, which might then bias results. The cosmological dependencies of the resulting HOD-informed EFT parameter distribution will be discussed in detail in Section 4.1.
@@
The dark matter halos are identified using the CompaSO halo finder, a highly efficient, on-the-fly group finder specifically designed for the AbacusSummit simulations [87]. We use the AbacusHOD code [88] to rapidly generate 10,000 power spectrum multipole measurements for each simulation box, corresponding to 10,000 different sets of HOD parameters. These measurements are produced using the HOD model described in Section 2.1, with the HOD parameters sampled according to the distributions listed in Table 1. The HOD parameter distributions are broad and based on the priors selected for the DESI LRG HOD study by [57], providing a conservative choice to inform the EFT parameter space. Unlike HOD fitting, our approach does not aim to imprint small-scale clustering information into the HOD-informed EFT parameter space. Instead, by using a broad range of HOD parameters and varying cosmologies, we ensure that the resulting HOD-informed EFT parameter distribution is primarily shaped by the structure of the HOD model and the selected HOD distributions for specific types of galaxies. It is important to note that the choice of HOD model and parameter distributions could impact the resulting EFT parameter space, as it directly reflects the characteristics of the galaxies being studied. However, we find that the results are insensitive to this choice, suggesting the primary impact of the HOD prior is to remove unphysical regions of EFT parameter space. A detailed discussion of the HOD parameter distribution choices will be presented in Section 4.2, highlighting how they shape the EFT parameter landscape.
@@ 3.2 Rapid best-fit estimation of EFT parameters
We fit the monopole and quadrupole up to $k_{\rm{max}}=0.2\,{\rm{Mpc}}^{-1}h$ as our baseline results, with the impact of including the hexadecapole discussed in Section 4.2. The choice of $k_{\rm{max}}=0.2\,{\rm{Mpc}}^{-1}h$ is reasonable given the samples and the volumes considered in the tests in Section 5; we thus match the same $k_{\rm{max}}$ in the best-fitting procedure When the hexadecapole is not included, the parameters $c_{\mathrm{r,2}}$ and $c_{\epsilon,1}$ are fixed to 0, as $c_{\mathrm{r,2}}$ is exactly degenerate with $c_{\mathrm{r,1}}$ and $c_{\epsilon,1}$ could not be well constrained by just the monopole and quadrupole. When the hexadecapole is included, we continue to use $k_{\rm{max}}=0.2\,{\rm{Mpc}}^{-1}h$ and allow all ten EFT parameters to vary. During the best-fit finding process, we fix the cosmology to match the corresponding simulation box and vary only the EFT parameters. Additionally, we set the model galaxy density to match the true galaxy density corresponding to the power spectrum we fit to, in order to consistently fit the stochastic coefficients across all samples. Given the broad range of the HOD parameter space, we use wide, uninformative flat priors for the best-fit estimation at this stage of the process. The broad priors used, along with the standard priors for EFTofLSS-based full-shape cosmological analysis, are listed in Table 2.
@@ 4.2 Influence of HOD sample distribution and hexadecapole inclusion
Figure 3: Contour plots illustrating the NF learned HOD-informed priors using three different configurations: TNHOD M+Q (red contours), UHOD M+Q (green contours), and TNHOD M+Q+H (blue contours). The unfilled gray contours (and boundary lines) represent the standard prior for EFTofLSS full-shape analysis as listed in Table 2 for comparison. Each contour indicates the 1-$\sigma$ and 2-$\sigma$ intervals of the distribution. The lower triangle compares the TNHOD M+Q and UHOD M+Q priors, highlighting the impact of removing the Gaussian component from the HOD sample distribution. The upper triangle compares the TNHOD M+Q and TNHOD M+Q+H priors, showing the effect of including the hexadecapole moment on the priors.
@@
When comparing all HOD-informed priors with the standard prior, it is clear that the HOD-informed priors, regardless of the configuration, provide provide a detailed picture of the degeneracy directions between the EFT parameters, which the standard prior fails to capture. Additionally, for most parameters, the 1-$\sigma$ contour areas are significantly smaller for the HOD-informed priors, indicating a noticeable reduction in the effective EFT parameter space. This reduction shows that the HOD-informed priors help guide the exploration of a more physically meaningful parameter space, mitigating prior effects and ultimately leading to more accurate cosmological parameter posteriors.
@@ 5 Comparative analysis of priors
AbacusSummit_base_c000_ph000 box at redshift $z=0.8$. The HOD parameters used are based on those described in table 3 of [57] for LRG with $0.80.6<z<$ 0.8, with the galaxy density tuned to $5\times 10^{-4}\,{\rm Mpc}^{-3}$. We computed the power spectrum multipoles for this catalog, adopting 18 linear bins with a scale cut of $0.02<k<0.2\,h^{-1}\rm{Mpc}$ for the monopole, quadrupole, and hexadecapole moments. For the covariance matrix, we utilized the Gaussian Covariance from CovaPT, with linear bias, volume, and number density matching the catalog.
@@
The results made using the standard prior are significantly biased with the true cosmological parameters far from the credible regions. This is evident when comparing the HOD-informed prior contours (red, green, and blue) with the standard prior contours (gray dashed, gray dot-dashed) of Figure 4. On the other hand, all HOD-informed priors provide credible regions more consistent with the underlying cosmology, particularly for parameters such as $\Omega_{c}h^{2}$, $\ln 10^{10}A_{s}$ and $n_{s}$ compared to the standard prior. For example, focusing on the TNHOD M+Q (red) contours, the $\sigma$-distance for $\Omega_{c}h^{2}$ improves from 1.65 with the standard prior to 0.57. Similarly, for $\ln 10^{10}A_{s}$, the value improves from 1.46 to 0.37, and for $n_{s}$ , it drops from 1.88 to 0.50. Also for $H_{0}$, there is a notable improvement, with the $\sigma$-distance reduced from 1.55 to 0.87. These comparisons clearly demonstrate the ability of the HOD-informed prior to mitigate the prior effect. Note that though power spectrum alone has limited constraining power on $n_{s}$, it is still correlated with other parameters and EFT parameters. The HOD-informed prior, by reducing the degeneracies among those parameters, allows $n_{s}$ to be better recovered. The HOD-informed prior leads to tighter constraints by $28\%$, $9\%$, $29\%$, and $37\%$ for $\Omega_{c}h^{2}$, $H_{0}$, $\ln 10^{10}A_{s}$, and $n_{s}$, respectively, compared to the standard prior. Thus we see that the $\sigma$-distance improvements were not driven by having larger credible intervals, but instead by mitigating prior effects. This improvement in constraints is not driven by physical knowledge of the particular halos being modelled, as we have used a wide range of HOD parameters.
@@
Another interesting point is that the posteriors using TNHOD and UHOD are nearly identical, indicating that the cosmological analysis, at least for the $\Lambda$CDM model, is not significantly influenced by the distribution of HOD sample used to generate the HOD-informed prior. Although the UHOD prior exhibits a more concentrated distribution, as shown in Figure 3, it does not lead to tighter constraints compared to the TNHOD prior.
@@
Figure 5 present the contour plots for the $w\mathrm{CDM}$ model. In this case, the fit using the standard prior suffers from more severe prior effects due to the expanded parameter space. The $\sigma$-distance values for most parameters are around or exceed 2 in the M+Q case. Although the inclusion of the hexadecapole in the M+Q+H case provides some improvement, there are still significant deviations, with $\Omega_{c}h^{2}$ and $n_{s}$ showing over 2-$\sigma$ differences from the true values. In contrast, the HOD-informed prior (red and blue contours) consistently recovers the true values across all parameters. Taking the TNHOD M+Q case as an example, the $\sigma$-distance for $\Omega_{c}h^{2}$ decreases from 1.90 with the standard prior to 1.03, for $H_{0}$ from 2.26 to 1.57, for $\ln 10^{10}A_{s}$ from 2.55 to 1.55, for $n_{s}$ from 1.97 to 1.13. and for the parameter $w$, from 1.97 to 1.42. The green (UHOD M+Q) and red (TNHOD M+Q) contours match each other perfectly, constraining power using HOD-informed prior remaining similar as using standard prior, and the arguments made for the $\Lambda$CDM case still hold for the $w$CDM model.

### v2

@@ 1 Introduction
In this paper, we use a sample of such scenarios to derive an HOD-informed prior for the EFTofLSS model to address the issue of an overly large nuisance parameter volume. Our method involves generating a diverse set of galaxy power spectrum measurements from HOD mock catalogs, covering a wide range of HOD parameter space and various cosmological models. By fitting the EFTofLSS model to the power spectra derived from these mock catalogs, we identify the best-fit EFTofLSS parameters and use their distribution to form our HOD-informed prior [59]. This approach ensures that the HOD-informed prior encompasses most of the physically plausible power spectra, allowing us to delineate a more constrained and physical parameter space. Regions outside this prior are likely unphysical and can be excluded from the analysis, thereby helping to reduce the prior effect and improve the precision of cosmological parameter estimates. We carefully tested several configurations that could influence the prior, including the impact of different cosmologies, variations in HOD parameter samples, the inclusion or exclusion of the hexadecapole moment, and dependence on covariance matrices. These tests ensure the robustness and reliability of the HOD-informed prior across various scenarios.
In [60, 61], a similar strategy was adopted by incorporating simulation-based priors into their full-shape analysis, calibrated via a field-level forward model. This approach reduced the prior volume for bias parameters and led to substantial improvements in constraints on $\Omega_{m}$ and $\sigma_{8}$. They also explored the cosmology dependence of their learned prior and found it to be mild. Our work leverages a novel EFTofLSS emulator, allowing us to stack fits from a broad range of cosmological models, thereby covering a wider physical volume and making our priors less sensitive to a cosmological baseline.
@@ 2.2 Galaxy power spectrum in the EFTofLSS
Our notation follows that established in [17], incorporating a combination of linear terms, 1-loop Standard Perturbation Theory (SPT) terms, counterterms, and stochastic terms. The quantity $\mu$ is the cosine of the angle between the LoS direction and wavenumber vector $\mathbf{k}$. $P_{11}(k)$ denotes the linear matter power spectrum, and $f$ is the growth factor. The scale $k_{\textsc{m}}^{-1}$ controls the spatial extension of the collapsed objects, the scale $k_{\textsc{r}}^{-1}$ controls the counterterms needed to define the product of velocities at the same point [78, 79], and $\bar{n}_{g}$ is the mean galaxy number density. In the following, we will set $k_{\textsc{m}}^{-1}=0.7\,h^{-1}\rm{Mpc}$ and $k_{\textsc{r}}^{-1}=0.35\,h^{-1}\rm{Mpc}$, while $\bar{n}_{g}$ is estimated for each box. $Z_{n}$ represents the $n$th order of the redshift galaxy density kernel,
@@ 3.1 Generating galaxy power spectra from HOD and N-body simulations
We use the AbacusSummit simulation suite [83, 84, 85, 86, 87] for this analysis, which provides a set of large, high-resolution N-body simulations specifically designed to meet the accuracy requirements of the DESI galaxy survey. These simulations cover a wide range of cosmological parameters, making them particularly suitable for our study. For this work, we utilize the AbacusSummit_base_c000_ph000 box, which represents the Planck 2018 cosmology [34], along with the AbacusSummit_base_c130_ph000 to AbacusSummit_base_c160_ph000 boxes at $z=0.8$. The latter form an unstructured emulator grid surrounding the c000 box, providing wider coverage of the 7-dimensional cosmological space $\{\omega_{b}h^{2},\omega_{c}h^{2},h,A_{s},n_{s},w_{0},w_{a}\}$. These 32 boxes were specifically selected to collectively sample variations in the cosmological parameters of interest, ensuring broad and scattered coverage across the parameter space. Each simulation box has a size of $2000\,h^{-1}\rm{Mpc}$ per side and contains $6912^{3}$ dark matter particles. This setup allows us to generate galaxy power spectra with diverse cosmological parameters, ensuring generality in our findings - the prior should not be tied to a single comsological model, which might then bias results. The cosmological dependencies of the resulting HOD-informed EFT parameter distribution will be discussed in detail in Section 4.1.
@@
The dark matter halos are identified using the CompaSO halo finder, a highly efficient, on-the-fly group finder specifically designed for the AbacusSummit simulations [88]. We use the AbacusHOD code [89] to rapidly generate 10,000 power spectrum multipole measurements for each simulation box, corresponding to 10,000 different sets of HOD parameters. These measurements are produced using the HOD model described in Section 2.1, with the HOD parameters sampled according to the distributions listed in Table 1. The HOD parameter distributions are broad and based on the priors selected for the DESI LRG HOD study by [57], providing a conservative choice to inform the EFT parameter space. Unlike HOD fitting, our approach does not aim to imprint small-scale clustering information into the HOD-informed EFT parameter space. Instead, by using a broad range of HOD parameters and varying cosmologies, we ensure that the resulting HOD-informed EFT parameter distribution is primarily shaped by the structure of the HOD model and the selected HOD distributions for specific types of galaxies. It is important to note that the choice of HOD model and parameter distributions could impact the resulting EFT parameter space, as it directly reflects the characteristics of the galaxies being studied. However, we find that the results are insensitive to the choice of parameter distribution, suggesting the primary impact of the HOD prior is to remove unphysical regions of EFT parameter space. A detailed discussion of the HOD parameter distribution choices will be presented in Section 4.2, highlighting how they shape the EFT parameter landscape. In future work, we will apply this method to the DESI data using multiple HOD models and parameter choices, further testing this dependence.
@@ 3.2 Rapid best-fit estimation of EFT parameters
We fit the monopole and quadrupole up to $k_{\rm{max}}=0.2\,{\rm{Mpc}}^{-1}h$ as our baseline results, with the impact of including the hexadecapole discussed in Section 4.2. The choice of $k_{\rm{max}}=0.2\,{\rm{Mpc}}^{-1}h$ is reasonable given the samples and the volumes considered in the tests in Section 5; we thus match the same $k_{\rm{max}}$ in the best-fitting procedure. When the hexadecapole is not included, the parameters $c_{\mathrm{r,2}}$ and $c_{\epsilon,1}$ are fixed to 0, as $c_{\mathrm{r,2}}$ is exactly degenerate with $c_{\mathrm{r,1}}$ and $c_{\epsilon,1}$ could not be well constrained by just the monopole and quadrupole. When the hexadecapole is included, we continue to use $k_{\rm{max}}=0.2\,{\rm{Mpc}}^{-1}h$ and allow all ten EFT parameters to vary. During the best-fit finding process, we fix the cosmology to match the corresponding simulation box and vary only the EFT parameters. Additionally, we set the model galaxy density to match the true galaxy density corresponding to the power spectrum we fit to, in order to consistently fit the stochastic coefficients across all samples. Given the broad range of the HOD parameter space, we use wide, uninformative flat priors for the best-fit estimation at this stage of the process. The broad priors used, along with the standard priors for EFTofLSS-based full-shape cosmological analysis, are listed in Table 2.
@@ 4.2 Influence of HOD sample distribution and hexadecapole inclusion
Figure 3: Contour plots illustrating the NF learned HOD-informed priors using three different configurations: TNHOD M+Q (red contours), UHOD M+Q (green contours), and TNHOD M+Q+H (blue contours). The unfilled gray contours (and boundary lines) represent the standard prior for EFTofLSS full-shape analysis as listed in Table 2 for comparison. Each contour indicates the 1-$\sigma$ and 2-$\sigma$ intervals of the distribution. The lower triangle compares the TNHOD M+Q and UHOD M+Q priors, highlighting the impact of removing the Gaussian component from the HOD sample distribution. Cross markers indicate the best-fit EFT parameters for the DESI-like Abacus mock power spectrum (monopole and quadrupole) used in Section 5, with cosmological parameters fixed to their true values. The upper triangle compares the TNHOD M+Q and TNHOD M+Q+H priors, showing the effect of including the hexadecapole moment on the priors.
@@
When comparing all HOD-informed priors with the standard prior, it is clear that the HOD-informed priors, regardless of the configuration, provide a detailed picture of the degeneracy directions between the EFT parameters, which the standard prior fails to capture. Additionally, for most parameters, the 1-$\sigma$ contour areas are significantly smaller for the HOD-informed priors, indicating a noticeable reduction in the effective EFT parameter space. This reduction shows that the HOD-informed priors help guide the exploration of a more physically meaningful parameter space, mitigating prior effects and ultimately leading to more accurate cosmological parameter posteriors. Since our power spectrum measurements are derived from finite-volume simulation boxes, the best-fit EFT parameters obtained from any single measurement will scatter around the “true” values that would be expected in an infinite-volume limit. This additional scatter, primarily driven by sample variance, effectively acts as a Gaussian smoothing on the HOD-informed priors, broadening them and making them more conservative. The priors still account for HOD variations and robustly capture the most plausible EFT parameter values, while conservatively excluding unphysical regions.
@@ 5 Comparative analysis of priors
AbacusSummit_base_c000_ph000 box at redshift $z=0.8$. The HOD parameters used are based on those described in table 3 of [57] for LRG with $0.80.6<z<$ 0.8, with the galaxy density tuned to $5\times 10^{-4}\,{\rm Mpc}^{-3}$. We computed the power spectrum multipoles for this catalog, adopting 18 linear bins with a scale cut of $0.02<k<0.2\,h^{-1}\rm{Mpc}$ for the monopole, quadrupole, and hexadecapole moments. In the lower triangle of Figure 3, the cross markers indicate the underlying best-fit EFT parameter values for the mock power spectrum monopole and quadrupole, obtained with cosmological parameters fixed to their true values. As expected, most of these EFT parameters lie near the centers of the HOD-informed priors, which are specifically tailored to DESI-like LRGs. For the covariance matrix, we utilized the Gaussian Covariance from CovaPT, with linear bias, volume, and number density matching the catalog.
@@
The results made using the standard prior are biased with the true cosmological parameters far from the credible regions. This is evident when comparing the HOD-informed prior contours (red, green, and blue) with the standard prior contours (gray dashed, gray dot-dashed) of Figure 4. On the other hand, all HOD-informed priors provide credible regions more consistent with the underlying cosmology, particularly for parameters such as $\Omega_{c}h^{2}$, $\ln 10^{10}A_{s}$ and $n_{s}$ compared to the standard prior. For example, focusing on the TNHOD M+Q (red) contours, the $\sigma$-distance for $\Omega_{c}h^{2}$ improves from 1.65 with the standard prior to 0.57. Similarly, for $\ln 10^{10}A_{s}$, the value improves from 1.46 to 0.37, and for $n_{s}$ , it drops from 1.88 to 0.50. Also for $H_{0}$, there is a notable improvement, with the $\sigma$-distance reduced from 1.55 to 0.87. These comparisons clearly demonstrate the ability of the HOD-informed prior to mitigate the prior effect. Note that though power spectrum alone has limited constraining power on $n_{s}$, it is still correlated with other parameters and EFT parameters. The HOD-informed prior, by reducing the degeneracies among those parameters, allows $n_{s}$ to be better recovered. The HOD-informed prior leads to tighter constraints by $28\%$, $9\%$, $29\%$, and $37\%$ for $\Omega_{c}h^{2}$, $H_{0}$, $\ln 10^{10}A_{s}$, and $n_{s}$, respectively, compared to the standard prior. Thus we see that the $\sigma$-distance improvements were not driven by having larger credible intervals, but instead by mitigating prior effects. This improvement in constraints is not driven by physical knowledge of the particular halos being modelled, as we have used a wide range of HOD parameters.
@@
Another interesting point is that the posteriors using TNHOD and UHOD are nearly identical, indicating that the cosmological analysis, at least for the $\Lambda$CDM model, is not significantly influenced by the distribution of HOD sample used to generate the HOD-informed prior. Although the UHOD prior exhibits a more concentrated distribution and the true $b_{1}$ lies beyond the 1-$\sigma$ interval of UHOD in the lower triangle of Figure 3, it does not produced biased or tighter constraints than the TNHOD prior.
@@
Figure 5 present the contour plots for the $w\mathrm{CDM}$ model. In this case, the fit using the standard prior suffers from more severe prior effects due to the expanded parameter space. The $\sigma$-distance values for most parameters are around or exceed 2 in the M+Q case. Although the inclusion of the hexadecapole in the M+Q+H case provides some improvement, there are still modest deviations, with $\Omega_{c}h^{2}$ and $n_{s}$ showing over 2-$\sigma$ differences from the true values. In contrast, the HOD-informed prior (red and blue contours) consistently recovers the true values across all parameters. Taking the TNHOD M+Q case as an example, the $\sigma$-distance for $\Omega_{c}h^{2}$ decreases from 1.90 with the standard prior to 1.03, for $H_{0}$ from 2.26 to 1.57, for $\ln 10^{10}A_{s}$ from 2.55 to 1.55, for $n_{s}$ from 1.97 to 1.13. and for the parameter $w$, from 1.97 to 1.42. The green (UHOD M+Q) and red (TNHOD M+Q) contours match each other perfectly, constraining power using HOD-informed prior remaining similar as using standard prior, and the arguments made for the $\Lambda$CDM case still hold for the $w$CDM model.